### v1

# AutoP2C: An LLM-Based Agent Framework for Code Repository Generation from Multimodal Content in Academic Papers

Abstract.
Machine Learning (ML) research is spread through academic papers featuring rich multimodal content, including text, diagrams, and tabular results. However, translating these multimodal elements into executable code remains a challenging and time-consuming process that requires substantial ML expertise. We introduce “Paper-to-Code” (P2C), a novel task that transforms the multimodal content of scientific publications into fully executable code repositories, which extends beyond the existing formulation of code generation that merely converts textual descriptions into isolated code snippets. To automate the P2C process, we propose AutoP2C, a multi-agent framework based on large language models that processes both textual and visual content from research papers to generate complete code repositories. Specifically, AutoP2C contains four stages: (1) repository blueprint extraction from established codebases, (2) multimodal content parsing that integrates information from text, equations, and figures, (3) hierarchical task decomposition for structured code generation, and (4) iterative feedback-driven debugging to ensure functionality and performance. Evaluation on a benchmark of eight research papers demonstrates the effectiveness of AutoP2C, which can successfully generate executable code repositories for all eight papers, while OpenAI-o1 or DeepSeek-R1 can only produce runnable code for one paper. The code is available at https://github.com/shoushouyu/Automated-Paper-to-Code.
Multimodality, Multi-agent framework, Large Language Model, Code Repository Generation
conference: Preprint; Under Review; ccs: Computing methodologies Multi-agent systemsccs: Human-centered computing Human computer interaction (HCI)
1. Introduction
Scientific advancements in the Machine Learning (ML) community are typically disseminated via academic papers (Hanson et al., 2024; Ayanwale et al., 2024; Li et al., 2022). Automating the conversion of these papers into code repositories accelerates both the reproducibility and the real-world application of scientific findings (Ball, 2023; Puangjaktha et al., 2024; Trofimova et al., 2024). As we know, these research papers generally encompass diverse multimodal content, such as textual descriptions, architectural diagrams, and tabular data (Song et al., 2025; Pramanick et al., 2024; Blecher et al., 2023). Multimodal content can enhance human comprehension and can also help generate corresponding code repositories. For example, visual diagrams help design complex model architectures, textual descriptions deepen the understanding of algorithmic implementations, and tabular hyperparameters can be used to construct configuration files.
Figure 1. Illustration of the “Paper-to-Code” (P2C) task. It involves processing multimodal content, such as text, images, and tables, from an academic paper to automatically generate a complete, executable code repository together with explanatory diagrams.
In this paper, we focus on how to automatically generate a code repository for an academic paper with multimodal content and call it the “Paper-to-Code” (P2C) task, as depicted in Fig. 1. Specifically, given a research paper containing heterogeneous, multimodal content, P2C aims to generate a complete executable code repository that accurately implements the described methods and reproduces the reported results. However, unlike traditional code generation, there are three novel challenges in dealing with the P2C task. First, P2C requires processing and integrating of multimodal inputs from academic papers (e.g., text, images, tables). Second, new model designs must be implemented accurately and align with the paper’s descriptions. Third, it needs to generate multi-file, structured code repositories rather than isolated code snippets (Sundar et al., 2024; Yang et al., 2023).
Existing approaches for automated ML implementation face various limitations when handling the P2C task. First, some related works involve traditional Automated Machine Learning (AutoML) methods (Ali and Smith, 2006; Bergstra and Bengio, 2012; Jin et al., 2019; Bisong and Bisong, 2019; Das et al., 2020), which are employed to automate the end-to-end process of training machine learning models for structured data. However, they rely on predefined workflows and model pools, rendering them unsuitable for implementing new architectures or unconventional algorithms. Second, Large Language Models (LLMs) have been applied to significantly enhance the flexibility of code generation from natural language in recent years (Liu et al., 2023; Mu et al., 2024). Nonetheless, on the one hand, even state-of-the-art models struggle to generate multi-file code repositories, constrained by context window limitations and inference complexity (Anil et al., 2022; Newman et al., 2020). On the other hand, these LLMs primarily process textual descriptions, neglecting the rich visual and structural multimodal information in research papers. Third, multi-agent frameworks (Yang et al., 2024; Tao et al., 2024) have emerged as a promising approach for complex code generation by employing specialized agents collaborating in iterative development cycles. For example, self-organized agents (Ishibashi and Nishimura, 2024) use a multi-agent system that automatically replicates agents based on task complexity, thereby enabling scalable large-scale code generation. However, these multi-agent methods predominantly focus on generic software development scenarios with text-only inputs, lacking the multimodal understanding required to interpret architectural diagrams, parse mathematical equations, and extract implementation details from tables and figures (Li et al., 2024). As a result, they fall short of providing in-depth research and practical capability for generating fully replicable code repositories from ML research papers.
To fill this research-to-implementation gap, we propose AutoP2C, a multi-agent framework specifically designed to generate complete code repositories from the multimodal content of research papers. AutoP2C comprises four stages. In the first stage, it extracts universal code structures from established repositories to construct architectural blueprints. The next stage involves multimodal content parsing, which integrates information from text, diagrams, and tables into a unified representation. In the third stage, divide-and-conquer task planning decomposes complex implementations into hierarchical subtasks with clearly defined interfaces. Finally, execution feedback-driven debugging localizes errors and aligns the code with the multimodal specifications of the paper through iterative testing. Unlike previous approaches that treat code generation as a unimodal text-to-text translation problem (Trofimova et al., 2024; Islam et al., 2024), AutoP2C takes advantage of multimodal understanding to capture the full spectrum of information presented in academic papers. To evaluate the performance of P2C, we propose two novel metrics to assess the structural completeness of code repositories and combine these with actual runtime performance to quantify how faithfully the generated code reproduces the paper’s reported performance.
The major contributions are threefold. (1) We introduce a novel task, i.e., P2C, and first formalize it as transforming the multimodal content of academic papers into structured, multifile code repositories with explanatory visualizations. (2) We propose AutoP2C, an innovative multi-agent framework that consists of four carefully designed stages to automate the P2C task. Furthermore, AutoP2C produces internal file dependency graphs that facilitate user understanding and review to improve transparency. (3) We establish a benchmark spanning multiple ML domains to evaluate P2C. Correspondingly, we propose two P2C evaluation metrics for structural completeness and combine them with run-time performance to serve as the metrics for the generated code repositories.
2. Related Works
2.1. Large Language Models for Code Generation
LLM have demonstrated their capabilities in code generation tasks, evolving from simple code completion to handling increasingly complex programming challenges (Nijkamp et al., 2022; Wang et al., 2023; Thakur et al., 2024; Le et al., 2022; Dong et al., 2024). Self-planning approaches (Jiang et al., 2024) have enhanced these LLMs by guiding them to outline solution steps based on user intent before generating code, resulting in more targeted implementations. Models like CodeGen (Nijkamp et al., 2022) have pioneered multistep program synthesis that improves the understanding of user requirements, while DeepSeek-Coder (Guo et al., 2024) employs a project-level code corpus with innovative “fill-in-the-blank” pre-training objectives to enhance code infill capabilities. Despite these advances, current LLMs for code generation operate only within a unimodal paradigm i.e., processing only textual input (Xu et al., 2024a; Anil et al., 2022; Zhang et al., 2023), which prevents them from effectively interpreting the rich multimodal content found in research papers, such as architectural diagrams, mathematical equations, and tabular results. Furthermore, these models typically generate isolated code snippets rather than cohesive, multi-file repositories that researchers require for implementation.
2.2. Multi-Agent for Code Generation
Single LLMs often struggle with context management, modular design, error detection, and complex reasoning in multiple files in code generation. Multi-agent frameworks address these shortcomings by distributing specialized tasks across collaborative agents that can maintain focused context awareness while collectively managing repository-level complexity (Huang et al., 2023; Wu et al., 2023; Chen et al., 2023; Qian et al., 2023). For example, CodeCoR (Pan et al., 2025) implements a self-reflective multi-agent where agents collaboratively generate code, develop test cases, and perform iterative repairs to improve both quality and executability. Similarly, MetaGPT (Hong et al., 2024) adopts a meta-programming paradigm that assigns distinct roles to specialized agents, enabling the coordinated production of coherent and maintainable multi-file code repositories through structured workflows. Among these frameworks, CodeAgent (Zhang et al., 2024) represents the first framework specifically designed for repository-level code generation. This LLM-based agent integrates five programming tools through four distinct agent strategies to manage complex code repositories effectively. Despite these notable advances, efficiently generating high-quality, multi-file code repositories remains challenging, particularly for implementing ML algorithms described in research papers as they require the understanding for the multimodal context.
Figure 2. Overview of the AutoP2C framework. It contains four stage, namely (1) repository blueprint extraction, which analyzes established ML repositories to generate a standardized blueprint for future code organization; (2) multimodal content parsing, which processes PDF papers into a unified, implementation-focused representation; (3) hierarchical task decomposition, which generates structured implementation plans; and (4) iterative feedback-driven implementation, which transforms task descriptions into an executable code repository.
3. Methods
3.1. Problem Formulation
The task of “Paper-to-Code” (P2C) is defined as the generation of executable code repositories based on multimodal content from academic papers on ML. Formally, we denote the multimodal content in the academic paper $P$ as a triple comprising three modalities, that is, $P\triangleq(\text{Texts},\text{Images},\text{Tables})$, where $\text{Texts}=\{\text{Paragraph}_{1},\ldots,\text{Paragraph}_{N}\}$ contains all textual content, such as abstract, method, and results; $\text{Images}=\{\text{Image}_{1},\ldots,\text{Image}_{A}\}$ represents visual elements, including architectural diagrams and results visualizations; and $\text{Tables}=\{\text{Table}_{1},\ldots,\text{Table}_{B}\}$ contains structured data, e.g., experimental results, hyperparameter configurations, and comparative analyzes presented in tabular format. The target output of P2C is a code repository $R\triangleq\big{(}\{\text{Module}_{1},\dots,\text{Module}_{k},\dots,\text{Module}_{K}\},\ D\big{)}$, where $\text{Module}_{k}$ represents a component of the implementation (e.g., data processor, model architecture, training procedures, evaluation metrics) and it actually is a collection of related source code files, i.e., $\text{Module}_{k}\triangleq\{\text{File}_{k,1},\text{File}_{k,2},\dots,\text{File}_{k,N_{k}}\}$. Here, $N_{k}\geq 1$ ensures that each $k$-th module contains at least one implementation file. Correspondingly, $D$ refers to the generated explanatory diagrams, such as dependency graphs, which further improve the understanding of the repository’s structure and facilitate navigation through the generated code repository. Based on the above definitions, the P2C task can be formalized as
(1) $f_{\,\text{P2C}}:P\rightarrow R.$
Importantly, this formulation distinguishes P2C from conventional code generation tasks (Lu et al., 2021; Austin et al., 2021; Gong et al., 2024; LaBash et al., 2024) in three aspects: (1) P2C processes multimodal context rather than purely textual descriptions for code generation; (2) P2C produces structured multi-file code repositories instead of isolated code snippets; (3) P2C generates explanatory multimodal visualizations that improve code comprehensibility.
Figure 3. Case study illustrates the complete workflow of AutoP2C for the EDoRA: … (Nasiri and Garraghan, 2025) paper. The process begins with (1) extracting multimodal content (text, images, tables) from the original research paper using MinerU (Wang et al., 2024), followed by (2) multimodal content parsing to produce refined, structured content with cross-modal relationships preserved. Based on this processed multimodal information, AutoP2C then performs (3) hierarchical task decomposition to generate detailed file descriptions with explicit references to paper sections. These descriptions guide (4) the implementation of functional code that accurately captures the paper’s algorithmic specifications, culminating in (5) the generation of explanatory diagrams that visualize component relationships in the resulting code repository.
3.2. Method Overview
To address the P2C task, we propose AutoP2C, a multi-agent framework based on LLM shown in Fig. 2. AutoP2C consists of four stages that imitate human experts implementing code repositories from research papers. The first stage is the repository blueprints extraction, where an LLM analyzes well-established ML code repositories of different domains =$\mathcal{R}=\{R_{1},R_{2},\dots,R_{m}\}$ to extract common architectural patterns as the “base knowledge”. Specifically, this stage produces a standardized repository organization template $T$ across four dimensions, namely repository architecture, file interdependencies, function designs, and class structures. This repository-level understanding provides scaffolding for subsequent code generation and is performed only once as a pre-processing step.
The second stage, multimodal content parsing, processes academic papers from PDF format to extract necessary information across modalities for code implementation. This stage employs the OCR model (Wei et al., 2024; Chen et al., 2025) for textual content extraction, the Vision Language Model (VLM) (Wu et al., 2024; Chu et al., 2024) for the understanding of architectural diagrams, the LLM for the analysis of mathematical equations for algorithmic formulations and the extraction of structured data for configurations and results. Afterward, an LLM synthesizes all these diverse elements from different modalities into a unified representation $P_{\text{distilled}}$ in text that preserves the major cross-modal relationships necessary for code implementation.
With the content of the distilled paper $P_{\text{distilled}}$ purely in text together with the repository blueprint $T$ at hand, the hierarchical task decomposition stage uses Large Reasoning Models (LRMs) (Guo et al., 2025) to generate a structured implementation plan with task descriptions $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$ for each component, defining file functionalities, dependencies, and interfaces while ensuring alignment with the paper’s algorithmic specifications. Concretely, the decomposition of AutoP2C follows a top-down design, beginning with high-level architectural decisions and progressively refining into specific implementation details.
The final stage, namely iterative feedback-driven implementation, transforms task descriptions $\mathcal{T}$ into a fully executable code repository $R_{\text{exec}}$ through an implementation-verification loop. Specifically, a coding agent generates initial implementations for each module based on their respective task description. Then, a verification agent executes the code, identifies errors or inconsistencies, and provides targeted feedback for refinement. This feedback drives subsequent refinement iterations until the repository successfully reproduces the results described in the paper. Additionally, at this stage, AutoP2C generates multimodal explanatory visualizations that enhance code comprehensibility by illustrating data flows throughout the generated code repository.
3.3. Repository Blueprint Extraction
To establish a repository organization framework that accommodates diverse implementation patterns, AutoP2C first derives standardized blueprints from existing high-quality ML code repositories before processing any specific paper. Given a set of representative code repositories =$\mathcal{R}=\{R_{1},R_{2},\dots,R_{m}\}$ that span various domains of ML, we employ LLMs to extract structural patterns in four hierarchical dimensions. The repositories are specifically selected based on three criteria: (1) popularity metrics in the ML community (e.g., $>1000$ GitHub stars), (2) code quality indicators (e.g., with comprehensive documentation), and (3) domain diversity to ensure coverage across computer vision, natural language processing, graph and others.
First, on the repository level, the LLM examines the overall architecture by quantifying folder organization patterns and file distribution statistics, denoted as $A_{\text{arch}}(R_{i})=\{(\text{folder}_{i}^{n},|F_{i}^{n}|,F_{i}^{n}=\{f^{n,1}_{i},f^{n,2}_{i},\dots,f^{n,j}_{i},\dots\})|R_{i}\in\mathcal{R},n\geq 1\}$, where $\text{folder}_{i}^{n}$ represents a top-level folder (e.g., “models”, “data”, “utils”) in $R_{i}$, $f_{i}^{n,j}$ denotes the individual files contained within that folder, and $|F_{i}^{n}|$ quantifies the number of files.
This analysis is performed through automated repository structure parsing combined with LLM-guided categorization that classifies folders according to their functional purpose. For instance, the LLM recognizes that folders containing predominantly class definitions represent model architectures, while those with data-loading operations represent preprocessing components.
Second, on the folder level, the LLM determines the relationships between folders by extracting the interdependencies of files in these folders and the workflow sequence for each file in repository $R_{i}$, denoted as $A_{\text{relationship}}(R_{i})=\{(f_{i}^{j},D_{i}^{j},O_{i}^{j})|R_{i}\in\mathcal{R},j\geq 1\}$, where for each file $f_{i}^{j}$, $D_{i}^{j}$ is its dependency graph encoded as a structured list of import statements and function calls to other modules, and $O_{i}^{j}$ captures its operational logic as a summarized text representation of the file’s primary functionality. Here, for simplicity, we ignore the folder information in $f_{i}^{j}$.
Third, on the function level, the LLM characterizes design patterns of functions in each file, denoted as $A_{\text{func\_design}}(R_{i})=\{(\text{func}_{i,j}^{a},\text{sig}_{i,j}^{a},\text{flow}_{i,j}^{a})|R_{i}\in\mathcal{R},a\geq 1\}$, where $\text{func}_{i,j}^{a}$ is the a-th function of the j-th file in repository $R_{i}$, represented by its signature $\text{sig}_{i,j}^{a}$ (including parameter types, default values, and return specifications) and execution flow $\text{flow}_{i,j}^{a}$ (a structured representation of control flow including conditional branches and loop structures).
Fourth, on the class level, the LLM identifies design patterns of classes that depict domain-specific concepts, denoted as $A_{\text{class\_design}}(R_{i})=\{(\text{class}_{i,j}^{b},\text{attr}_{i,j}^{b},\text{method}_{i,j}^{b})|R_{i}\in\mathcal{R},b\geq 1\}$, where $\text{class}_{i,j}^{b}$ is the b-th class of the j-th file in repository $R_{i}$, characterized by its attributes (i.e., $\text{attr}_{i,j}^{b}$), which define the class’s state, such as configuration parameters, model weights, or dataset properties, and methods (i.e., $\text{method}_{i,j}^{b}$) that define the class’s behavior, including initialization, forward passes, and utility functions.
After analyzing individual repositories, the repository blueprint extraction stage synthesizes a unified template by identifying common patterns across repositories,
(2) $T=$ $\text{LLM}_{\text{synthesize}}(\{A_{\text{arch}}(R_{i}),A_{\text{relationship}}(R_{i}),$
$A_{\text{func\_design}}(R_{i}),A_{\text{class\_design}}(R_{i})\}_{i=1}^{m}).$
The synthesize operation employs LLM to identify recurring structures across multiple repositories. This process involves frequency analysis to determine common directory structures (e.g., data/models/utils pattern), abstraction of implementation details to create generalized function templates, and identification of standard interface conventions between components. The resulting template $T$ is represented as a hierarchical markdown structure with four main sections that correspond to the above analyzed dimensions.
3.4. Multimodal Content Parsing
The multimodal parsing process begins with transforming the PDF format of paper $P$ into a markdown format by using an OCR model MinerU (Wang et al., 2024). This initial conversion yields a raw representation $P_{\text{raw}}=\text{OCR}_{\text{MinerU}}(P)=(T_{\text{raw}},I_{\text{raw}},M_{\text{raw}},D_{\text{raw}})$, where $T_{\text{raw}}$ represents textual content that includes methodology descriptions and algorithmic procedures; $I_{\text{raw}}$ contains visual elements such as architectural diagrams, flow charts, and result visualizations; $M_{\text{raw}}$ contains mathematical equations; and $D_{\text{raw}}$ is tabular data such as hyperparameter configurations and experimental results. However, this raw conversion typically introduces artifacts such as fragmented paragraphs, misplaced image references, and misformed equations that must be addressed before future content extraction (Liu et al., 2024). To restore structural integrity, AutoP2C employs an LLM to identify and correct these conversion artifacts, i.e. $T_{\text{structured}}=\text{LLM}_{\text{restore}}(T_{\text{raw}})$.
Next, AutoP2C tries to integrate multimodal content. For visual content processing, AutoP2C employs VLMs (Wu et al., 2024; Chu et al., 2024) to extract semantic understanding from visual elements to generate structured descriptions that capture component relationships, depicted as $I_{\text{parsed}}=\{\text{Parse}_{\text{visual}}(i)|i\in I_{\text{raw}}\}$. Rather than relying on VLMs to directly describe the images, we design a guiding prompt consisting of three key components: (1) Emphasizing only code-relevant information by disregarding superfluous visual details (e.g., color), so as to avoid introducing noise and distracting attention. (2) Providing a comprehensive description, particularly including all numerical elements in the image to prevent omission. (3) Leveraging the image’s caption to establish strong correlations between the image and relevant text from the paper, thereby producing an accurate, text-aligned answer to minimize ambiguity. Similarly, mathematical equations are translated from typeset representations into computational forms via LLM i.e., $M_{\text{parsed}}=\{\text{Parse}_{\text{math}}(m)|m\in M_{\text{raw}}\}$. Tabular data undergoes similar extraction to identify hyperparameters, experimental configurations, and evaluation metrics by $D_{\text{parsed}}=\{\text{Parse}_{\text{tabular}}(d)|d\in D_{\text{raw}}\}$.
Finally, after processing individual modalities, AutoP2C integrates with LLM to remove redundancy, namely $P_{\text{integrated}}=\text{LLM}_{\text{integrate}}(T_{\text{structured}},I_{\text{parsed}},M_{\text{parsed}},D_{\text{parsed}})$. The final stage of multimodal parsing filters the integrated content to focus on code implementation-relevant information with LLM, namely $P_{\text{distilled}}=\text{LLM}_{\text{filter}}(P_{\text{integrated}})$. This filtering process identifies and retains content directly relevant to code implementation, including algorithmic descriptions, architectural specifications, hyperparameter settings, and evaluation methodologies, while deprioritizing content such as literature reviews, theoretical justifications, and references that contribute less directly to implementation.
3.5. Hierarchical Task Decomposition
With the distilled multimodal content $P_{\text{distilled}}$ and the repository blueprint $T$, AutoP2C progresses to decomposing the implementation into hierarchically structured tasks with LLM. The decomposition process follows a top-down design that progressively refines implementation requirements across multiple levels of abstraction. First, AutoP2C generates a repository-level architecture that defines the overall structure with LLM i.e., $A_{\text{repo}}=\text{LLM}_{\text{architecture}}(P_{\text{distilled}},T)=\{(F_{i},\phi_{i},S_{i})\}_{i=1}^{n}$, where $F_{i}$ indicates a file in the repository and is associated with its description of the functionality $\phi_{i}$ and the corresponding source sections $S_{i}$ of the academic paper $P_{\text{distilled}}$. It uses the repository blueprint $T$ to ensure that the generated architecture adheres to established best practices while accommodating the unique contributions of the article.
Second, AutoP2C generates detailed component specifications for each file in $A_{\text{repo}}$, namely $C_{i}=\text{LLM}_{\text{components}}(F_{i},P_{\text{distilled}})=\{(\kappa_{j},\alpha_{j},\mu_{j})\}_{j=1}^{m_{i}}$, where $\kappa_{j}$ represents a class or function, $\alpha_{j}$ defines its attributes or parameters, $\mu_{j}$ specifies its methods or behaviors, and $m_{i}$ is the number of classes or functions in $F_{i}$. To prevent semantic drift (Ji et al., 2023) i.e, the gradual deviation from the original algorithmic intent during implementation, each component is explicitly linked to specific sections of the paper that describe its algorithmic or architectural details. Meanwhile, AutoP2C constructs explicit modeling of inter-file dependencies, namely $D=\text{LLM}_{\text{MapDependencies}}({C_{i}}_{i=1}^{n})=\{(F_{i},\{(F_{j},\epsilon_{i,j})\})\}$ where $\epsilon_{i,j}$ represents the specific components from file $F_{j}$ upon which file $F_{i}$ depends. This dependency mapping ensures structural coherence across the code repository and facilitates proper interface design between components. Importantly, AutoP2C prioritizes internal dependencies over external library dependencies.
Third, to mitigate the complexity of reasoning for LLMs when designing intricate implementation details, AutoP2C employs an iterative refinement approach to generate task descriptions for each file, denoted as $\tau_{i}=\text{LLM}_{\text{Gen}}(F_{i},C_{i},D,P_{\text{distilled}})$. This process focuses on one file at a time, refining its implementation details while maintaining awareness of the global repository structure. For each file, AutoP2C produces step-by-step implementation guidelines that explicitly refer to relevant sections of the paper containing algorithmic specifications, mathematical formulations, or architectural diagrams. Finally, the set of task descriptions $\mathcal{T}=\{\tau_{1},\tau_{2},...,\tau_{n}\}$ serves as a blueprint for implementation.
Additionally, to enhance interpretability and facilitate human review, AutoP2C also generates a visual representation of inter-file dependencies. This visualization is created through a graph rendering process that transforms the formal dependency structure $D$ into an interactive diagram. The nodes in the graph represent files, while the directed edges represent dependencies, with edge labels indicating the specific nature of each dependency (e.g., inheritance, import). The visualization uses color coding to differentiate between component types (e.g., models, training).
3.6. Iterative Feedback-Driven Implementation
The final stage of AutoP2C transforms hierarchical task descriptions $\mathcal{T}=\{\tau_{1},\tau_{2},...,\tau_{n}\}$ into a fully executable code repository through an iterative implementation-verification process. Specifically, AutoP2C employs a progressive file-by-file implementation approach
(3) $Code_{i}=\text{LLM}_{\text{ImplementCode}}(\tau_{i},\{Code_{j}|j<i\},P_{\text{distilled}}),$
where each code file $Code_{i}$ is generated based on its corresponding task description $\tau_{i}$, previously implemented code files $\{Code_{j}|j<i\}$, and the distilled paper representation $P_{\text{distilled}}$. This sequential approach ensures that the code implementation remains consistent across the repository, with each new file building on the existing implementation context. The implementation order follows the dependency graph established during the hierarchical task decomposition stage, beginning with foundational components and progressively implementing dependent modules.
Next, AutoP2C performs a validation process to ensure the fidelity of implemented code files to the core algorithms of the paper i.e., $V=\text{LLM}_{\text{validate}}(\{Code_{i}\}_{i=1}^{n},P_{\text{distilled}})$. This validation focuses on three aspects: (1) model architecture design, verifying that neural network structures match the paper’s specifications; (2) loss function formulation, ensuring mathematical correctness in the implementation of objective functions; and (3) optimization strategy, confirming that update rules align with the described training procedures. The validation process employs cross-modal verification, comparing code implementations with relevant sections of the paper through textual descriptions, architectural diagrams, and mathematical formulations to ensure complete alignment.
Despite the validation step, producing a fully executable pipeline in a single pass remains challenging due to complex inter-file interactions and interface consistency requirements. To address this, AutoP2C implements an iterative debugging mechanism with LLM, namely $\{Code_{i}^{\prime}\}_{i=1}^{n}=\text{LLM}_{\text{Debug}}(\{Code_{i}\}_{i=1}^{n},E)$, where $\{Code_{i}^{\prime}\}_{i=1}^{n}$ represents the refined code and $E$ denotes execution errors encountered during testing. The debugging process operates through a two-phase approach. First, the error localization phase, i.e., $L=\text{LLM}_{\text{LocalizeError}}(E,\{Code_{i}\}_{i=1}^{n})$, identifies the specific files and components associated with each LLM execution error, minimizing the scope of necessary modifications and reducing the risk of unintended side effects. Second, the error correction phase, i.e., $\{Code_{i}^{\prime}\}=\text{LLM}_{\text{CorrectError}}(L,\{Code_{i}\},P_{\text{distilled}})$, generates targeted modifications to resolve identified issues while maintaining adherence to the paper’s specifications, which may involve adjusting function interfaces, aligning data structures, or refining implementation details to ensure compatibility across components.
After successful debugging, AutoP2C integrates hyperparameter optimization (HPO) through an automated process following previous work (Luo et al., 2024). Specifically, we incorporate ray.tune (Raiaan et al., 2024) into the main execution file, enabling exploration of the hyperparameter space defined in the paper. The HPO implementation extracts relevant hyperparameters from the experimental sections of the paper and configures appropriate search spaces based on reported values and ranges.
Table 1. Comparative evaluation of AutoP2C against baseline models (OpenAI-o1 and DeepSeek-R1) across eight research papers in terms of four metrics, i.e., absolute performance (accuracy on target datasets), relative performance (ratio to original implementation), class completeness ($\text{COMP}_{\text{class}}$) and function completeness ($\text{COMP}_{\text{func}}$). Red crosses ($\times$) indicate failure to generate executable code. The rightmost columns show computational resource consumption through input and output token counts for the AutoP2C framework. Results demonstrate AutoP2C’s superior ability to translate multimodal paper content into functional code repositories, with improvements in both execution success and structural completeness.
Paper’s Title ArXiv ID Task Type Performance Relative Performance $\text{COMP}_{\text{class}}$ $\text{COMP}_{\text{func}}$ $\text{Tokens}_{\text{in}}$ $\text{Tokens}_{\text{out}}$
o1 R1 ours o1 R1 ours o1 R1 ours o1 R1 ours
Distilled… (Fontana et al., 2024) None multiclass $\times$ $\times$ 82.8% $\times$ $\times$ 89.8% 90.8% 90.8% 90.8% 31.3% 37.8% 60.2% 852K 120K
EDoRA: … (Nasiri and Garraghan, 2025) 2501.12067 binary $\times$ $\times$ 82.5% $\times$ $\times$ 89.9% 27.1% 0% 54.3% 42.5% 17.0% 88.9% 458K 93K
No More Adam:… (Xu et al., 2024b) 2412.11768 multiclass 81.8% 75.3% 91.4% 87.5% 80.5% 97.8% 43.1% 24.8% 86.1% None None None 1177K 145K
Entropy Aware… (Nazari et al., 2024) 2403.04636 multiclass $\times$ $\times$ 79.9% $\times$ $\times$ 103.0% 15.3% 3.0% 53.8% 9.1% 8.0% 46.2% 334K 77K
$L^{2}$:… (Liang et al., 2024) 2403.06064 multiclass $\times$ $\times$ 79.5% $\times$ $\times$ 96.5% 11.0% 50.9% 75.9% 18.4% 12.5% 31.4% 477K 106K
ReffAKD:… (Doshi and Kim, 2024) 2404.09886 multiclass $\times$ $\times$ 91.6% $\times$ $\times$ 98.0% 13.0% 11.7% 25.9% 23.8% 23.2% 35.9% 547K 112K
KD-LoRA:… (Azimi et al., 2024) 2410.20777 binary $\times$ $\times$ 91.1% $\times$ $\times$ 99.0% None None None 60.0% 20.0% 80.0% 310K 88K
Convolutional… (Linse et al., 2023) 2411.18388 multiclass $\times$ $\times$ 92.0% $\times$ $\times$ 122.0% 43.7% 36.7% 73.1% 18.5% 4.4% 18.5% 501K 98K
4. Experiments
4.1. Evaluation Metrics
In our experiments, we evaluate the functional correctness and structural completeness of the code repositories generated by the AutoP2C framework. For functional correctness, we execute the generated code repositories and the code of the original papers on the same datasets. To compare their performance, we introduce two traditional performance metrics, namely (1) absolute performance, which measures the outcomes (e.g., accuracy, F1 score, or other task-specific metrics) achieved by the AutoP2C generated code on designated datasets; and (2) relative performance, calculated as the ratio between the performance of the AutoP2C generated code and that of the original implementation. To evaluate structural completeness, i.e., how accurately the code structure captures the multimodal context in the academic paper, we define two new LLM-based metrics based on the LLM-as-a-Judge paradigm (Zheng et al., 2023; He et al., 2025), leveraging the LLM’s outstanding performance in code quality evaluation. The first is function completeness ($\text{COMP}_{\text{func}}$), which quantifies the quality of the implementation of the functions using a weighted LLM-based scoring, depicted as
(4) $\text{COMP}_{\text{func}}=\frac{\sum_{func_{i}\in Func}w_{i}\cdot\text{score}_{\text{func}}(f_{i})}{\sum_{func_{i}\in Func}w_{i}},$
where $Func$ represents the set of functions in the code of the original paper, $w_{i}$ denotes the number of lines in function $func_{i}$, and $\text{score}_{\text{func}}(f_{i})\in\{0,0.2,0.4,0.6,0.8,1\}$ reflects implementation completeness on a 6-score scale given by LLM. The scoring ranges from unimplemented (00) to fully implemented ($1$), with intermediate values representing progressive levels of implementation quality: correct purpose but flawed logic ($0.2$), correct logic with major omissions ($0.4$), correct logic with partial omissions ($0.6$), and minor omissions ($0.8$). Similarly, class completeness ($\text{COMP}_{\text{class}}$) evaluates how faithfully the generated code implements class structures specified in the paper’s diagrams and text:
(5) $\text{COMP}_{\text{class}}=\frac{\sum_{c\in C}\sum_{m\in M_{c}}w_{m}\cdot\text{score}_{\text{method}}(m)}{\sum_{c\in C}\sum_{m\in M_{c}}w_{m}},$
where $C$ represents the set of classes, $M_{c}$ denotes the methods in class $c$, $w_{m}$ indicates the code lines of the method $m$, and $\text{score}_{\text{method}}(m)$ follows the same 6-point scoring system as $\text{score}_{\text{func}}$. Notably, if the implementations enhance functionality beyond the paper’s specifications while maintaining alignment with its core principles, it will receive a $+0.2$ bonus, capped at $1.0$. Additionally, we track computational resource utilization by measuring the number of input and output tokens throughout the entire process, which provides insights into the framework’s efficiency when processing multimodal content with different complexities.
4.2. Benchmark and Baseline
The P2C task represents a novel challenge that extends beyond traditional code generation by requiring the integration of multimodal content into coherent, multi-file code repositories. Given the absence of specialized benchmarks or frameworks designed specifically for this task, we first establish a new benchmark comprising eight recent research papers all published after 2024 (Fontana et al., 2024; Nasiri and Garraghan, 2025; Nazari et al., 2024; Xu et al., 2024b; Liang et al., 2024; Doshi and Kim, 2024; Azimi et al., 2024; Linse et al., 2023). These papers are selected from paperswithcode.com to evaluate the capability of AutoP2C in processing diverse modal datasets across various ML domains. The benchmark spans six ML tasks (including training strategy optimization (Xu et al., 2024b), node classification (Nazari et al., 2024; Liang et al., 2024), model compression (Doshi and Kim, 2024), parameter-efficient fine-tuning (Azimi et al., 2024; Nasiri and Garraghan, 2025), network pruning (Fontana et al., 2024), and image classification (Linse et al., 2023)) and three modal types of datasets (containing computer vision (Fontana et al., 2024; Xu et al., 2024b; Doshi and Kim, 2024; Linse et al., 2023), natural language processing (Nasiri and Garraghan, 2025; Azimi et al., 2024), and graph (Nazari et al., 2024; Liang et al., 2024) related datasets). We also established baselines using two state-of-the-art LLMs: OpenAI’s recently released o1 (Zhong et al., 2024) and DeepSeek-R1 (Guo et al., 2025), both known for their advanced code generation capabilities.
4.3. Experiment Settings
From the repository blueprint extraction stage to the hierarchical task decomposition stage, we employ GPT-4o due to its advanced multimodal understanding capability. For the initial file implementation, where code generation from processed specifications is the primary focus, we utilize o1-mini, a model optimized for code generation tasks. Then, we use o1 in the validation process, which is good at detecting nuanced logical inconsistencies. Iterative code refinement is handled by o3-mini, which excels at refinement tasks. To ensure reproducibility and consistency across experiments, we standardize the decoding temperature to 0 for all temperature-configurable LLMs. All experiments are conducted on an Intel(R) Xeon(R) Platinum 8352V CPU and an NVIDIA A40 (48GB) GPU.
Table 2. Ablation study demonstrating the impact of each AutoP2C component on code generation performance. It examines four components, namely, repository blueprint extraction (blueprint), multimodal content parsing (MM), hierarchical task decomposition (Decop.), and iterative feedback-driven implementation (Feedback). Each row shows results when removing one component (marked with ‘x’) while keeping others (marked with checkmarks). The bottom row shows the complete AutoP2C framework with all components enabled. Performance metrics include absolute accuracy (Perf.), relative performance compared to original implementations (Rel. Perf.), and structural fidelity metrics ($\text{COMP}_{\text{class}}$ and $\text{COMP}_{\text{func}}$). Red crosses ($\times$) indicate failure to generate executable code.
Modules Paper’s Title: Distilled… (Fontana et al., 2024) Paper’s Title: ReffAKD:… (Doshi and Kim, 2024) Paper’s Title: $L^{2}$:… (Liang et al., 2024) Paper’s Title: Entropy Aware… (Nazari et al., 2024)
Blueprint MM Decomp. Feedback Perf. Rel. Perf. $\text{COMP}_{\text{class}}$ $\text{COMP}_{\text{func}}$ Perf. Rel. Perf. $\text{COMP}_{\text{class}}$ $\text{COMP}_{\text{func}}$ Perf. Rel. Perf. $\text{COMP}_{\text{class}}$ $\text{COMP}_{\text{func}}$ Perf. Rel. Perf. $\text{COMP}_{\text{class}}$ $\text{COMP}_{\text{func}}$
x 10.0% 10.8% 90.8% 25.0% 90.3% 96.6% 16.3% 32.0% 66.9% 81.2% 16.9% 18.1% 38.4% 49.5% 4.2% 17.9%
x 10.2% 11.0% 90.8% 44.1% 83.5% 89.4% 16.0% 35.9% 70.2% 85.2% 0.0% 19.3% 50.1% 64.5% 45.1% 46.2%
x 66.2% 71.8% 90.8% 23.5% 10.2% 10.9% 8.4% 32.5% 13.0% 15.8% 7.4% 9.6% 14.0% 18.0% 16.6% 13.6%
x $\times$ $\times$ 13.7% 24.6% $\times$ $\times$ 16.1% 24.3% $\times$ $\times$ 0.0% 16.7% $\times$ $\times$ 33.6% 34.9%
82.8% 89.8% 90.8% 60.2% 91.6% 98.0% 25.9% 35.9% 79.5% 96.5% 75.9% 31.4% 79.9% 103.0% 53.8% 46.2%
4.4. Results
4.4.1. Evaluation for Functional Correctness
The results in Table 1 demonstrate the advantages of AutoP2C in both implementation feasibility and functional fidelity compared to the baselines. While OpenAI-o1 and DeepSeek-R1 successfully generate executable code for only one paper (“No More Adam: …” (Xu et al., 2024b) ), AutoP2C produces functional implementations across all eight research papers in our benchmark. This contrast in success rate (100% versus 12.5%) underscores the effectiveness of AutoP2C in performing the P2C task. Even in the single case where the baseline models produce executable code, AutoP2C achieves superior performance. Specifically, for “No More Adam: …” (Xu et al., 2024b), our AutoP2C delivers 91.4% accuracy compared to 81.8% for o1 (9.6% improvement) and 75.3% for R1 (16.1% improvement). The relative performance metric further validates the effectiveness of our framework, with AutoP2C achieving between 89.8% and 122.0% of the performance of the original implementations in all eight papers. Notably, in two cases, the performance of the code repositories generated by AutoP2C exceeds that of the code of the original papers, namely “Entropy Aware..” (Nazari et al., 2024) (103.0%) and “Convolutional…” (Linse et al., 2023) (122.0%). The consistent high relative performance (averaging 99.5% across all papers) confirms that AutoP2C’s staged generation paradigm ensures both compliance with paper specifications and operational integrity in real-world execution. Additionally, we provide a case study of AutoP2C in Fig. 3, demonstrating how our approach successfully transforms multimodal research content into executable code through its four-stage pipeline.
4.4.2. Evaluation for Structural Completeness
The structural completeness metrics in Table 1 provide quantitative evidence of AutoP2C’s superior ability. For class completeness ($\text{COMP}_{\text{class}}$), AutoP2C consistently outperforms both baseline models across all papers where class implementations were relevant. Our framework achieves an average score of 65.7%, representing a 1.9×1.9$\times$1.9 × increase compared to OpenAI-o1 (34.9%) and a 2.1×2.1$\times$2.1 × increase compared to DeepSeek-R1 (31.1%), which demonstrates the effectiveness of AutoP2C in correctly interpreting architectural diagrams and translating them into appropriate class hierarchies and relationships. Similarly, for function completeness ($\text{COMP}_{\text{func}}$), AutoP2C delivers superior results in all applicable cases. On average, AutoP2C achieves a score of 51.6%, surpassing o1 (29.1%) by 1.8×1.8$\times$1.8 × and R1 (17.6%) by 2.9×2.9$\times$2.9 ×. The performance gap between AutoP2C and the baselines becomes even more pronounced when handling papers with complex mathematical notation. For graph-related papers specifically (Nazari et al., 2024; Liang et al., 2024), which typically contain extensive mathematical formulations, AutoP2C’s $\text{COMP}_{\text{class}}$ score is 4.9×4.9$\times$4.9 × higher than o1 and 2.4×2.4$\times$2.4 × higher than R1. Similarly, its $\text{COMP}_{\text{func}}$ score for these papers is 2.8×2.8$\times$2.8 × higher than o1 and 3.8×3.8$\times$3.8 × higher than R1. These substantial differences highlight AutoP2C’s particular strength in processing mathematically intensive content that requires careful integration of multiple modalities.
Table 3. Ablation study on the contribution of different modalities in multimodal content phrasing stage, including architectural diagrams (image) and experimental results (table) when processing the “Convolutional…” (Liang et al., 2024) paper. Checkmarks indicate included modalities, while ‘x’ denotes excluded ones.
Modality Paper’s Title: Convolutional…
Text Image Table Perf. Rel. Perf. $\text{COMP}_{\text{class}}$ $\text{COMP}_{\text{func}}$
x 70.1% 92.9% 59.8% 18.5%
x 88.9% 117.9% 36.7% 7.6%
92.0% 122.0% 73.1% 18.5%
4.5. Ablation Study
The ablation analysis in Table 2 reveals that the removal of any component degrades performance. When the repository blueprint extraction component is removed, the LLM must construct file organization structures without guidance, resulting in disorganized and redundant code. This absence of a structured template causes average performance to decline from 83.5% to 51.4%. The structural completeness metrics also decline, with the average class completeness dropping from 61.6% to 32.1% and the completeness of functions decreasing from 43.4% to 23.2%. The multimodal content parsing component is equally important. In particular, the absence of visual understanding and redundancy removing results in average performance declining to 53.5%, with class completeness and function completeness falling to 38.0% and 36.4% respectively. Eliminating the hierarchical task decomposition phase forces the model to generate entire implementations in a single pass, which results in the average class completeness decreasing to 30.8% and produces the lowest average function completeness (19.8%) scores across our ablation. Without the iterative feedback-driven implementation, AutoP2C relies entirely on single-pass generation, resulting in completely non-executable implementations across all test cases and yielding the lowest average class completeness of 15.9%. Moreover, the function completeness falls to 25.1%.
Table 3 presents another ablation study on the contribution of different modalities in the multimodal content parsing stage. The results demonstrate that each modality provides important information for generating accurate and complete code repositories. When architectural diagrams (image modality) are excluded from the input, performance decreases from 92.0% to 70.1% (a 21.9% absolute reduction), while class completeness declines from 73.1% to 59.8%. This substantial degradation occurs because architectural diagrams contain essential visual information about model structure, component relationships, and data flow that cannot be fully captured by textual descriptions alone. The absence of tabular modality produces different but equally important effects. These findings confirm the need for multimodal understanding in the P2C task.
5. Conclusion
This paper introduces “Paper-to-Code” (P2C), a novel task that addresses the challenging problem of translating multimodal research content from academic papers into fully executable code repositories. To tackle this task, we propose AutoP2C, a multi-agent framework that employs LLMs, which represents an advancement toward improving research reproducibility. Future work could explore the extension of AutoP2C to support additional programming languages beyond Python and incorporate user feedback to continuously improve the generated code repositories.